# On tori periods of Weil representations of unitary groups

### v1

@@ 4.3.4. Non-vanishing of global theta lifts for one-dimensional spaces over an étale algebra
Let $\mathbf{E}$ be an étale algebra of degree $n$ over $\mathbf{F}$, such that $L_{\mathbf{E}}^{1}$ is anisotropic, and let $\lambda\in\mathbf{E}^{\times}$. Choose a trace zero element $\delta\in\mathbf{K}^{\times}$. Consider the same setup as in Section 4.2. The non-vanishing result described in Section 4.3.3 has a straightforward extension that allows us to determine whether the global theta lift of an automorphic character of $\operatorname{U}(\mathrm{W}_{\mathbf{F},1}^{\delta})\left(\mathbb{A}_{\mathbf{F}}\right)$ to $\operatorname{U}(\mathrm{V}_{\mathbf{E},\lambda})\left(\mathbb{A}_{\mathbf{F}}\right)$ is non-zero. This extension will be important for our main result. Let $i_{\mathrm{V}}^{\prime}\colon\operatorname{U}(\mathrm{V}_{\mathbf{E},\lambda})\to L_{\mathbf{E}}^{1}$ and $i_{\mathrm{W}}^{\prime}\colon\operatorname{U}(\mathrm{W}_{\mathbf{E},1}^{\delta})\to L_{\mathbf{E}}^{1}$ be admissible isomorphisms. Let $\alpha\colon\left[L_{\mathbf{E}}^{1}\right]\to\mathbb{C}^{\times}$ be an automorphic character.
@@
  • For every place $v$, the big theta lift $\Theta_{{L_{\mathbf{E},\lambda,v}},{\mathrm{W}_{\mathbf{E},1,v}^{\delta}},\tilde{\iota}_{v},\psi_{v}}\left(\alpha_{v}\circ i^{\prime}_{\mathrm{V},v}\right)$ does not vanish.
  • The central $L$-function value $\mathcal{L}\left(\chi^{-1}_{\mathrm{W}_{\mathbf{F},1}^{\delta}}\circ N_{L_{\mathbf{E}}/\mathbf{K}}\cdot\alpha_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}\right)$ is non-zero.
@@ 5.1. Local problem
Let $F$ be a local field (either archimedean or non-archimedean of characteristic $\neq 2$) and let $K/F$ be a quadratic étale algebra. Let $\mathrm{V}$ be a non-degenerate $n$-dimensional hermitian space over $K$, and let $\mathrm{W}$ be a non-degenerate one-dimensional skew-hermitian space over $K$. Let $i_{\mathrm{W}}^{\prime}\colon\operatorname{U}(\mathrm{W})\rightarrow K^{1}$ be an admissible isomorphism (it is unique). Fix a character $\beta\colon K^{1}\to\mathbb{C}^{\times}$, and let $\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)=\Theta_{\mathrm{W},\mathrm{V},\tilde{\iota},\psi}(\beta\circ i^{\prime}_{\mathrm{W}})$ be the big theta lift of $\beta\circ i^{\prime}_{\mathrm{W}}$ to $\operatorname{U}(\mathrm{V})$, where $\tilde{\iota}$ is the splitting associated with the characters $\left(\chi_{\mathrm{W}},\chi_{\mathrm{V}}\right)$. Given a maximal torus $\mathrm{T}\subset\operatorname{U}(\mathrm{V})$ and a character $\alpha^{\prime}\colon\mathrm{T}\rightarrow\mathbb{C}^{\times}$, we would like to investigate whether the space $\operatorname{Hom}_{\mathrm{T}}\left(\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right),\alpha^{\prime}\right)$ is non-zero. By Theorem 2.3, we have that if $\mathrm{T}\subset\operatorname{U}(\mathrm{V})$ is a maximal torus, then there exists an $n$-dimensional étale algebra $E$ over $F$, an element $\lambda\in E^{\times}$, and an isomorphism $r\colon\mathrm{V}_{E,\lambda}\rightarrow\mathrm{V}$ of hermitian spaces, such that
@@
Let $E$ be an étale algebra of degree $n$ over $F$, $\alpha\colon L_{E}^{1}\rightarrow\mathbb{C}^{\times}$ be a character, and let $i\colon L_{E}^{1}\to\operatorname{U}(\mathrm{V})$ be an admissible embedding, corresponding to the data $\lambda\in E^{\times}$ and $r\colon\mathrm{V}_{E,\lambda}\to\mathrm{V}$. Then $\operatorname{Hom}_{i\left(L_{E}^{1}\right)}\left(\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right),\alpha\circ i^{-1}\right)$ is non-zero if and only if
@@
Henceforth, we will identify $\mathrm{W}$ with $\mathrm{W}_{F,1}^{\delta}$. Let us be in the setup of Section 3.2. We will use the following seesaw diagram:
@@
where $\operatorname{U}(\mathrm{W})$ is embedded in $\operatorname{U}(\mathrm{W}_{E,1}^{\delta})$ diagonally, i.e., $\operatorname{U}(\mathrm{W})$ acts on elements of $\mathrm{W}_{E,1}^{\delta}$ by scalar multiplication via the admissible isomorphism $i^{\prime}_{\mathrm{W}}\colon\operatorname{U}(\mathrm{W})\rightarrow K^{1}$. By the local seesaw identity (6), we have that
@@
where $i^{\prime}_{\mathrm{W}_{E,1}^{\delta}}\colon\operatorname{U}(\mathrm{W}_{E,1}^{\delta})\to L_{E}^{1}$ is an admissible isomorphism. Since $i^{\prime}_{\mathrm{W}_{E,1}^{\delta}}$ is an admissible isomorphism, we have that ${i^{\prime}_{\mathrm{W}_{E,1}^{\delta}}}_{\restriction_{\operatorname{U}(\mathrm{W})}}$ agrees with $i^{\prime}_{\mathrm{W}}$. Since the theta lift $\Theta\left(\alpha\circ i^{-1}\right)$ is a character, given that it is not zero, we have that the space $\operatorname{Hom}_{\operatorname{U}(\mathrm{W})}\left(\Theta\left(\alpha\circ i^{-1}\right)_{\restriction_{\operatorname{U}(\mathrm{W})}},\beta\circ i^{\prime}_{\mathrm{W}}\right)$ is non-zero if and only if $\Theta\left(\alpha\circ i^{-1}\right)\restriction_{\operatorname{U}(\mathrm{W})}$ is the same as $\beta\circ i^{\prime}_{\mathrm{W}}$. By (9), this is equivalent to
@@
as required. Finally, if $\operatorname{Hom}_{\operatorname{U}(\mathrm{W})}\left(\Theta\left(\alpha\circ i^{-1}\right)\restriction_{\operatorname{U}(\mathrm{W})},\beta\circ i^{\prime}_{\mathrm{W}}\right)$ is non-zero, it has to be one-dimensional, since all the representations involved are characters. ∎
@@
If $F$ is non-archimedean and $K/F$ is a quadratic field extension, then there exist exactly two isomorphism classes of non-degenerate hermitian spaces over $K$ of dimension $n$. The isomorphism class of such hermitian space is determined by its discriminant. We may use this to determine when there exists an admissible embedding $i\colon L_{E}^{1}\to\operatorname{U}(\mathrm{V})$ with non-zero $\operatorname{Hom}$-space. Let $\lambda\in E^{\times}$ be such that
@@
Then there exists an admissible embedding $i\colon L_{E}^{1}\to\operatorname{U}(\mathrm{V})$, such that
$\operatorname{Hom}_{i\left(L_{E}^{1}\right)}\left(\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right),\alpha\circ i^{-1}\right)\neq 0,$
if and only if $\beta=\left(\chi_{\mathrm{W}}^{-n}\cdot\chi_{\mathrm{V}}\right)\circ j_{F}^{-1}\cdot\alpha_{\restriction_{K^{1}}}$ and
$\operatorname{disc}\mathrm{V}=N_{E/F}\left(\lambda\right)\cdot\operatorname{disc}_{F}\left(E\right).$
@@
Let $E$ be an étale algebra of degree $n$ over $F$, and let $\alpha\colon L_{E}^{1}\rightarrow\mathbb{C}^{\times}$ be a character. Then we have that
@@
where the sum over $\mathrm{H}$ is over a set of representatives of isomorphism classes of non-degenerate hermitian spaces of degree $n$, and the sum over $i$ is over a set of representatives for $\Sigma_{E,\mathrm{H}}$. Here, $\Theta_{\mathrm{W},\mathrm{H}}\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)$ is the big theta lift from $\operatorname{U}(\mathrm{W})$ to $\operatorname{U}(\mathrm{H})$, with respect to a splitting corresponding to a prescribed choice of characters $\left(\chi_{\mathrm{V}},\chi_{\mathrm{W}}\right)$
@@
where the summation is over a set of representatives $\lambda\in E^{\times}$ for $E^{\times}/N_{L_{E}/E}\left(L_{E}^{\times}\right)$, where $\mathrm{H}_{\lambda}$ is the representative for the class of the hermitian space $\mathrm{V}_{E,\lambda}$, and where $i_{r_{\lambda}}\colon L_{E}^{1}\rightarrow\mathrm{T}_{E,r_{\lambda}}\subset\operatorname{U}(\mathrm{H}_{\lambda})$ is an arbitrary admissible embedding corresponding to the data $r_{\lambda}\colon\mathrm{V}_{E,\lambda}\rightarrow\mathrm{H}_{\lambda}$. Given such $\lambda$, we have by Theorem 5.1 that the space $\operatorname{Hom}_{i_{r_{\lambda}}(L_{E}^{1})}\left(\Theta_{\mathrm{W},\mathrm{H}_{\lambda}}\left(\beta\circ i^{\prime}_{\mathrm{W}}\right),\alpha\circ i_{r_{\lambda}}^{-1}\right)$ can be non-zero only when $\beta=\left(\chi_{\mathrm{W}}^{-n}\cdot\chi_{\mathrm{V}}\right)\circ j_{F}^{-1}\cdot\alpha\restriction_{K^{1}}$, and in this case the space is non-zero only for one class in $E^{\times}/N_{L_{E}/E}\left(L_{E}^{\times}\right)$, and for that class it is one-dimensional. Therefore, we get the result. ∎
@@ 5.2. Global problem
Let $\mathbf{F}$ be a number field, and let $\mathbf{K}/\mathbf{F}$ be a quadratic field extension. Let $\mathrm{V}$ be a non-degenerate $n$-dimensional hermitian space over $\mathbf{K}$, and let $\mathrm{W}$ be a non-degenerate one-dimensional skew-hermitian space over $\mathbf{K}$. Let $i^{\prime}_{\mathrm{W}}\colon\operatorname{U}(\mathrm{W})\to L_{\mathbf{F}}^{1}$ be an admissible isomorphism. Fix an automorphic character $\beta\colon\left[L_{\mathbf{F}}^{1}\right]\to\mathbb{C}^{\times}$ and let $\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)=\Theta_{\mathrm{W},\mathrm{V},\tilde{\iota},\psi}\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)$ be the global theta lift of $\beta\circ i^{\prime}_{\mathrm{W}}$ to $\operatorname{U}(\mathrm{V})\left(\mathbb{A}_{\mathbf{F}}\right)$, where $\tilde{\iota}$ is the splitting associated with the characters $\left(\chi_{\mathrm{V}},\chi_{\mathrm{W}}\right)$. Given a maximal anisotropic torus $\mathrm{T}\subset\operatorname{U}(\mathrm{V})$ and an automorphic character $\alpha\colon\left[\mathrm{T}\right]\to\mathbb{C}^{\times}$, we would like to investigate whether the $\alpha$-period of $\mathrm{T}$ is identically zero on $\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)$, that is, we would like to check whether the integral
@@
is zero for every $f\in\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)$. As before, given such $\mathrm{T}$, by Theorem 2.3, we may find an étale algebra $\mathbf{E}$ of degree $n$ over $\mathbf{F}$, an element $\lambda\in\mathbf{E}^{\times}$, and an isomorphism $r\colon\mathrm{V}_{E,\lambda}\to\mathrm{V}$ of hermitian spaces over $\mathbf{K}$, such that
@@
If $\mathcal{P}_{i\left(L_{\mathbf{E}}^{1}\right),\alpha\circ i^{-1}}$ is non-zero then for every $v$, we must have that the space $\operatorname{Hom}_{i\left(L_{\mathbf{E}}\right)\left(\mathbf{F}_{v}\right)}\left(\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)_{v},\alpha_{v}\circ i^{-1}_{v}\right)$ is non-zero. By Theorem 5.1 and the compatibility between the global and local theta lifts, this is equivalent to the first two conditions.
@@
As in Section 3.2.1, $\operatorname{U}(\mathrm{W})$ is realized as a subgroup of $\operatorname{U}(\mathrm{W}_{\mathbf{E},1}^{\delta})$ diagonally. By the results of Section 4.2, we have that
@@
where $\varphi\in\mathcal{S}\left(\mathrm{Y},\mathbb{A}_{\mathbf{F}}\right)$. Hence, the period $\mathcal{P}_{i\left(L_{\mathbf{E}}^{1}\right),\alpha\circ i^{-1}}$ is non-zero if and only if the global theta lift $\Theta_{\tilde{\iota}}\left(\alpha\circ i^{-1}\right)$ from $\operatorname{U}(L_{\mathbf{E},\lambda})\left(\mathbb{A}_{\mathbf{F}}\right)$ to $\operatorname{U}(\mathrm{W}_{\mathbf{E},1}^{\delta})\left(\mathbb{A}_{\mathbf{F}}\right)$ is non-zero, and the integral (10) is non-zero. By Theorems 4.3 and 3.5, the global theta lift $\Theta_{\tilde{\iota}}\left(\alpha\right)$ is non-zero if and only if conditions (2) and (3) hold. In this case, we have that
@@
where $i^{\prime}_{\mathrm{W}_{\mathbf{E},1}^{\delta}}\colon\operatorname{U}(\mathrm{W}_{\mathbf{E},1}^{\delta})\to L_{\mathbf{E}}^{1}$ is an admissible isomorphism. Hence, by choosing $\varphi$ such that $\theta_{\tilde{\iota}}\left(\varphi,\alpha\right)\neq 0$, we get from the fact that two different characters of a group are orthogonal, that if $\Theta_{\tilde{\iota}}\left(\alpha\circ i^{-1}\right)$ is not zero, then $\mathcal{P}_{i\left(L_{\mathbf{E}}^{1}\right),\alpha\circ i^{-1}}$ is not identically zero if and only if
@@
Since $i_{\mathrm{W}}^{\prime}$ and $i^{\prime}_{\mathrm{W}_{\mathbf{E},1}^{\delta}}$ are admissible isomorphisms, we have that they agree on $\operatorname{U}(\mathrm{W}_{\mathbf{F},1}^{\delta})\left(\mathbb{A}_{\mathbf{F}}\right)$, and therefore the condition is equivalent to
@@
Moreover, when these conditions hold, the isomorphism class of $\mathrm{H}$ as a hermitian space, and the class $\left[i\right]\in\Sigma_{\mathbf{E},\mathrm{H}}$ are unique. Here, $\Theta_{\mathrm{W},\mathrm{H}}\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)$ is the global theta lift of $\beta\circ i^{\prime}_{\mathrm{W}}$ from $\operatorname{U}(\mathrm{W})\left(\mathbb{A}_{\mathbf{F}}\right)$ to $\operatorname{U}(\mathrm{H})\left(\mathbb{A}_{\mathbf{F}}\right)$, taken with respect to the splitting defined by a prescribed choice of characters $\left(\chi_{\mathrm{V}},\chi_{\mathrm{W}}\right)$.

### v2

@@ 4.3.4. Non-vanishing of global theta lifts for one-dimensional spaces over an étale algebra
Let $\mathbf{E}$ be an étale algebra of degree $n$ over $\mathbf{F}$, such that $L_{\mathbf{E}}^{1}$ is anisotropic, and let $\lambda\in\mathbf{E}^{\times}$. Choose a trace zero element $\delta\in\mathbf{K}^{\times}$. Consider the same setup as in Section 4.2. The non-vanishing result described in Section 4.3.3 has a straightforward extension that allows us to determine whether the global theta lift of an automorphic character of $\operatorname{U}(\mathrm{W}_{\mathbf{F},1}^{\delta})\left(\mathbb{A}_{\mathbf{F}}\right)$ to $\operatorname{U}(\mathrm{V}_{\mathbf{E},\lambda})\left(\mathbb{A}_{\mathbf{F}}\right)$ is non-zero. This extension will be important for our main result.
Let $i_{\mathrm{V}}^{\prime}\colon\operatorname{U}(\mathrm{V}_{\mathbf{E},\lambda})\to L_{\mathbf{E}}^{1}$ and $i_{\mathrm{W}}^{\prime}\colon\operatorname{U}(\mathrm{W}_{\mathbf{E},1}^{\delta})\to L_{\mathbf{E}}^{1}$ be the obvious isomorphisms. Let $\alpha\colon\left[L_{\mathbf{E}}^{1}\right]\to\mathbb{C}^{\times}$ be an automorphic character.
@@
  • For every place $v$, the big theta lift
    $\Theta_{{L_{\mathbf{E},\lambda,v}},{\mathrm{W}_{\mathbf{E},1,v}^{\delta}},\tilde{\iota}_{v},\psi_{v}}\left(\alpha_{v}\circ i^{\prime}_{\mathrm{V},v}\right)$
    does not vanish.
  • The central $L$-function value
    $\mathcal{L}\left(\chi^{-1}_{\mathrm{W}_{\mathbf{F},1}^{\delta}}\circ N_{L_{\mathbf{E}}/\mathbf{K}}\cdot\alpha_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}\right)$
    is non-zero.
@@ 5.1. Local problem
Let $F$ be a local field (either archimedean or non-archimedean of characteristic $\neq 2$) and let $K/F$ be a quadratic étale algebra. Let $\mathrm{V}$ be a non-degenerate $n$-dimensional hermitian space over $K$, and let $\mathrm{W}$ be a non-degenerate one-dimensional skew-hermitian space over $K$. Let $i_{\mathrm{W}}^{\prime}\colon\operatorname{U}(\mathrm{W})\to K^{1}$ be the obvious isomorphism. Fix a character $\beta\colon K^{1}\to\mathbb{C}^{\times}$, and let $\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)=\Theta_{\mathrm{W},\mathrm{V},\tilde{\iota},\psi}(\beta\circ i^{\prime}_{\mathrm{W}})$ be the big theta lift of $\beta\circ i^{\prime}_{\mathrm{W}}$ to $\operatorname{U}(\mathrm{V})$, where $\tilde{\iota}$ is the splitting associated with the characters $\left(\chi_{\mathrm{W}},\chi_{\mathrm{V}}\right)$. Recall that in this case $\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)$ coincides with the small theta lift $\theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)=\theta_{\mathrm{W},\mathrm{V},\tilde{\iota},\psi}(\beta\circ i^{\prime}_{\mathrm{W}})$ because $\beta\circ i^{\prime}_{\mathrm{W}}$ is supercuspidal.
Given a maximal torus $\mathrm{T}\subset\operatorname{U}(\mathrm{V})$ and a character $\alpha^{\prime}\colon\mathrm{T}\to\mathbb{C}^{\times}$, we would like to investigate whether the space $\operatorname{Hom}_{\mathrm{T}}\left(\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right),\alpha^{\prime}\right)$ is non-zero.
By Theorem 2.3, we have that if $\mathrm{T}\subset\operatorname{U}(\mathrm{V})$ is a maximal torus, then there exists an $n$-dimensional étale algebra $E$ over $F$, an element $\lambda\in E^{\times}$, and an isomorphism $r\colon\mathrm{V}_{E,\lambda}\to\mathrm{V}$ of hermitian spaces, such that
@@
Let $E$ be an étale algebra of degree $n$ over $F$, $\alpha\colon L_{E}^{1}\to\mathbb{C}^{\times}$ be a character, and let $i\colon L_{E}^{1}\to\operatorname{U}(\mathrm{V})$ be an admissible embedding, corresponding to the data $\lambda\in E^{\times}$ and $r\colon\mathrm{V}_{E,\lambda}\to\mathrm{V}$. Then $\operatorname{Hom}_{i\left(L_{E}^{1}\right)}\left(\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right),\alpha\circ i^{-1}\right)$ is non-zero if and only if
@@
Henceforth, we will identify $\mathrm{W}$ with $\mathrm{W}_{F,1}^{\delta}$.
Let us be in the setup of Section 3.2. We will use the following seesaw diagram:
@@
where $\operatorname{U}(\mathrm{W})$ is embedded in $\operatorname{U}(\mathrm{W}_{E,1}^{\delta})$ diagonally, i.e., $\operatorname{U}(\mathrm{W})$ acts on elements of $\mathrm{W}_{E,1}^{\delta}$ by scalar multiplication via the obvious isomorphism $i^{\prime}_{\mathrm{W}}\colon\operatorname{U}(\mathrm{W})\to K^{1}$.
By the local seesaw identity (6), we have that
@@
where $i^{\prime}_{\mathrm{W}_{E,1}^{\delta}}\colon\operatorname{U}(\mathrm{W}_{E,1}^{\delta})\to L_{E}^{1}$ is the obvious isomorphism. Since $i^{\prime}_{\mathrm{W}_{E,1}^{\delta}}$ is the obvious isomorphism, we have that ${i^{\prime}_{\mathrm{W}_{E,1}^{\delta}}}_{\restriction_{\operatorname{U}(\mathrm{W})}}$ agrees with $i^{\prime}_{\mathrm{W}}$.
Since the theta lift $\Theta\left(\alpha\circ i^{-1}\right)$ is a character, given that it is not zero, we have that the space $\operatorname{Hom}_{\operatorname{U}(\mathrm{W})}\left(\Theta\left(\alpha\circ i^{-1}\right)_{\restriction_{\operatorname{U}(\mathrm{W})}},\beta\circ i^{\prime}_{\mathrm{W}}\right)$ is non-zero if and only if $\Theta\left(\alpha\circ i^{-1}\right)\restriction_{\operatorname{U}(\mathrm{W})}$ is the same as $\beta\circ i^{\prime}_{\mathrm{W}}$. By (9), this is equivalent to
@@
as required.
Finally, if $\operatorname{Hom}_{\operatorname{U}(\mathrm{W})}\left(\Theta\left(\alpha\circ i^{-1}\right)\restriction_{\operatorname{U}(\mathrm{W})},\beta\circ i^{\prime}_{\mathrm{W}}\right)$ is non-zero, it has to be one-dimensional, since all the representations involved are characters. ∎
@@
By substituting $\mathrm{W}=\mathrm{W}_{F,1}^{\delta}$ and $\left(\chi_{\mathrm{W}},\chi_{\mathrm{V}}\right)=\left(\mu,\mu^{n}\right)$ as in Section 3.1.2, we obtain Theorem 1.2.
Remark 5.3.
If $F$ is non-archimedean and $K/F$ is a quadratic field extension, then there exist exactly two isomorphism classes of non-degenerate hermitian spaces over $K$ of dimension $n$. The isomorphism class of such hermitian space is determined by its discriminant. We may use this to determine when there exists an admissible embedding $i\colon L_{E}^{1}\to\operatorname{U}(\mathrm{V})$ with non-zero $\operatorname{Hom}$-space, where $E$ is an arbitrary étale algebra of degree $n$ over $F$.
Let $\alpha\colon L_{E}^{1}\to\mathbb{C}^{\times}$ be a character and let $\lambda\in E^{\times}$ be such that
@@
(By Theorem 5.1, this is the only possible class $\lambda\in E^{\times}/N_{L_{E}/E}\left(L_{E}^{\times}\right)$, such that there exists an admissible embedding corresponding to $\lambda$ with non-vanishing Hom-space).
Then there exists an admissible embedding $i\colon L_{E}^{1}\to\operatorname{U}(\mathrm{V})$ corresponding to an isomorphism $r\colon\mathrm{V}_{E,\lambda}\to\mathrm{V}$ if and only if
$\operatorname{disc}\mathrm{V}=N_{E/F}\left(\lambda\right)\cdot\operatorname{disc}_{F}\left(E\right).$ (10)
In this case,
$\operatorname{Hom}_{i\left(L_{E}^{1}\right)}\left(\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right),\alpha\circ i^{-1}\right)\neq 0,$ (11)
if and only if
$\beta=\left(\chi_{\mathrm{W}}^{-n}\cdot\chi_{\mathrm{V}}\right)\circ j_{F}^{-1}\cdot\alpha_{\restriction_{K^{1}}}$ (12)
@@
If $F=\mathbb{R}$ and $K=\mathbb{C}$, then (10) is only a necessary condition for the existence of an isomorphism $r\colon\mathrm{V}_{E,\lambda}\to\mathrm{V}$. Given that such isomorphism exists, we have by Theorem 5.1 that (11) holds if and only (12) holds.
@@
Let $E$ be an étale algebra of degree $n$ over $F$, and let $\alpha\colon L_{E}^{1}\to\mathbb{C}^{\times}$ be a character. Then we have that
@@
where the sum over $\mathrm{H}$ is over a set of representatives of isomorphism classes of non-degenerate hermitian spaces of degree $n$, and the sum over $i$ is over a set of representatives for $\Sigma_{E,\mathrm{H}}$. Here, $\Theta_{\mathrm{W},\mathrm{H}}\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)$ is the big theta lift from $\operatorname{U}(\mathrm{W})$ to $\operatorname{U}(\mathrm{H})$, with respect to a splitting corresponding to a prescribed choice of characters $\left(\chi_{\mathrm{W}},\chi_{\mathrm{V}}\right)$.
@@
where the summation is over a set of representatives $\lambda\in E^{\times}$ for $E^{\times}/N_{L_{E}/E}\left(L_{E}^{\times}\right)$, where $\mathrm{H}_{\lambda}$ is the representative for the class of the hermitian space $\mathrm{V}_{E,\lambda}$, and where $i_{r_{\lambda}}\colon L_{E}^{1}\to\mathrm{T}_{E,r_{\lambda}}\subset\operatorname{U}(\mathrm{H}_{\lambda})$ is an arbitrary admissible embedding corresponding to the data $r_{\lambda}\colon\mathrm{V}_{E,\lambda}\rightarrow\mathrm{H}_{\lambda}$. Given such $\lambda$, we have by Theorem 5.1 that the space $\operatorname{Hom}_{i_{r_{\lambda}}(L_{E}^{1})}\left(\Theta_{\mathrm{W},\mathrm{H}_{\lambda}}\left(\beta\circ i^{\prime}_{\mathrm{W}}\right),\alpha\circ i_{r_{\lambda}}^{-1}\right)$ can be non-zero only when $\beta=\left(\chi_{\mathrm{W}}^{-n}\cdot\chi_{\mathrm{V}}\right)\circ j_{F}^{-1}\cdot\alpha\restriction_{K^{1}}$, and in this case the space is non-zero only for one class in $E^{\times}/N_{L_{E}/E}\left(L_{E}^{\times}\right)$, and for that class it is one-dimensional. Therefore, we get the result. ∎
Remark 5.5.
By substituting the same data as in Remark 5.2, we get Theorem 1.3.
@@ 5.2. Global problem
Let $\mathbf{F}$ be a number field, and let $\mathbf{K}/\mathbf{F}$ be a quadratic field extension. Let $\mathrm{V}$ be a non-degenerate $n$-dimensional hermitian space over $\mathbf{K}$, and let $\mathrm{W}$ be a non-degenerate one-dimensional skew-hermitian space over $\mathbf{K}$. Let $i^{\prime}_{\mathrm{W}}\colon\operatorname{U}(\mathrm{W})\to L_{\mathbf{F}}^{1}$ be the obvious isomorphism. Fix an automorphic character $\beta\colon\left[L_{\mathbf{F}}^{1}\right]\to\mathbb{C}^{\times}$ and let $\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)=\Theta_{\mathrm{W},\mathrm{V},\tilde{\iota},\psi}\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)$ be the global theta lift of $\beta\circ i^{\prime}_{\mathrm{W}}$ to $\operatorname{U}(\mathrm{V})\left(\mathbb{A}_{\mathbf{F}}\right)$, where $\tilde{\iota}$ is the splitting associated with the characters $\left(\chi_{\mathrm{W}},\chi_{\mathrm{V}}\right)$.
Given a maximal anisotropic torus $\mathrm{T}\subset\operatorname{U}(\mathrm{V})$ and an automorphic character $\alpha\colon\left[\mathrm{T}\right]\to\mathbb{C}^{\times}$, we would like to investigate whether the $\alpha$-period of $\mathrm{T}$ is identically zero on $\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)$, that is, we would like to check whether the integral
@@
is zero for every $f\in\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)$.
As before, given such $\mathrm{T}$, by Theorem 2.3, we may find an étale algebra $\mathbf{E}$ of degree $n$ over $\mathbf{F}$, an element $\lambda\in\mathbf{E}^{\times}$, and an isomorphism $r\colon\mathrm{V}_{E,\lambda}\to\mathrm{V}$ of hermitian spaces over $\mathbf{K}$, such that
@@
If $\mathcal{P}_{i\left(L_{\mathbf{E}}^{1}\right),\alpha\circ i^{-1}}$ is non-zero then for every $v$, we must have that
$\operatorname{Hom}_{i\left(L_{\mathbf{E}}\right)\left(\mathbf{F}_{v}\right)}\left(\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)_{v},\alpha_{v}\circ i^{-1}_{v}\right)\neq 0.$ (13)
If $\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)$ lies in the space of square-integrable automorphic forms, then by Theorem 5.1 and the compatibility between the global and local theta lifts, (13) is equivalent to the first two conditions of the theorem. We mention that if $\dim\mathrm{V}\neq 2$, then $\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)$ is square-integrable. When $\mathrm{V}$ is one-dimensional this follows from the fact that the groups are anisotropic, and when $\dim\mathrm{V}\geq 3$, this follows from [35, Corollary 10.1 part (4)] (in the notations of [35], $\rho_{n}=1$, $\dim\mathrm{V}=m\geq 3$ and $0\leq r\leq\left[\frac{m}{2}\right]$ so $m-r>1=\rho_{n}$).
@@
As in Section 3.2.1, $\operatorname{U}(\mathrm{W})$ is realized as a subgroup of $\operatorname{U}(\mathrm{W}_{\mathbf{E},1}^{\delta})$ diagonally.
By the results of Section 4.2, we have that
@@
where $\varphi\in\mathcal{S}\left(\mathrm{Y},\mathbb{A}_{\mathbf{F}}\right)$. Hence, the period $\mathcal{P}_{i\left(L_{\mathbf{E}}^{1}\right),\alpha\circ i^{-1}}$ is non-zero if and only if the global theta lift $\Theta_{\tilde{\iota}}\left(\alpha\circ i^{-1}\right)$ from $\operatorname{U}(L_{\mathbf{E},\lambda})\left(\mathbb{A}_{\mathbf{F}}\right)$ to $\operatorname{U}(\mathrm{W}_{\mathbf{E},1}^{\delta})\left(\mathbb{A}_{\mathbf{F}}\right)$ is non-zero, and the integral (14) is non-zero. By Theorems 4.4 and 3.6, the global theta lift $\Theta_{\tilde{\iota}}\left(\alpha\right)$ is non-zero if and only if conditions (2) and (3) hold. In this case, we have that
@@
where $i^{\prime}_{\mathrm{W}_{\mathbf{E},1}^{\delta}}\colon\operatorname{U}(\mathrm{W}_{\mathbf{E},1}^{\delta})\to L_{\mathbf{E}}^{1}$ is the obvious isomorphism. Hence, by choosing $\varphi$ such that $\theta_{\tilde{\iota}}\left(\varphi,\alpha\right)\neq 0$, we get from the fact that two different characters of a group are orthogonal, that if $\Theta_{\tilde{\iota}}\left(\alpha\circ i^{-1}\right)$ is not zero, then $\mathcal{P}_{i\left(L_{\mathbf{E}}^{1}\right),\alpha\circ i^{-1}}$ is not identically zero if and only if
@@
Since $i_{\mathrm{W}}^{\prime}$ and $i^{\prime}_{\mathrm{W}_{\mathbf{E},1}^{\delta}}$ are the obvious isomorphisms, we have that they agree on $\operatorname{U}(\mathrm{W}_{\mathbf{F},1}^{\delta})\left(\mathbb{A}_{\mathbf{F}}\right)$, and therefore the condition is equivalent to
@@
Moreover, when these conditions hold, the isomorphism class of $\mathrm{H}$ as a hermitian space, and the class $\left[i\right]\in\Sigma_{\mathbf{E},\mathrm{H}}$ are unique.
Here, $\Theta_{\mathrm{W},\mathrm{H}}\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)$ is the global theta lift of $\beta\circ i^{\prime}_{\mathrm{W}}$ from $\operatorname{U}(\mathrm{W})\left(\mathbb{A}_{\mathbf{F}}\right)$ to $\operatorname{U}(\mathrm{H})\left(\mathbb{A}_{\mathbf{F}}\right)$, taken with respect to the splitting defined by a prescribed choice of characters $\left(\chi_{\mathrm{W}},\chi_{\mathrm{V}}\right)$.